OpenStudy (anonymous):

Are the ratios 25/45 and 15/27 proportional? Explain. Help Please!

OpenStudy (anonymous):

You have to make sure if there's any number that you can divide them by numbers and get both of them to that number. For making sure you need to do this: a/b = c/d → a * d = b * d and see if the equation (proportion) is correct or not. (What i did up there is multiply both of the ratios by the numbers sitting in the bottom of each ratio) So now: 25/45 = 15/27 → 25 * 27 = ? = 15 * 45 → 675 = ? = 675 which is correct so the ratios ARE proportional!

OpenStudy (anonymous):